OpenStudy (anonymous):

A company will need $55,000 in 8 years for a new addition. To meet this goal, the company his to deposit money in an account today that pay 5% annual interest compoun quartely. Find the amount that should be invested to total $55,000 in 8 years.

OpenStudy (chris):

Check this out - http://www.algebralab.org/Word/Word.aspx?file=Algebra_InterestII.xml

OpenStudy (anonymous):

I am not getting it, please help me!!!!!!!!!

OpenStudy (chris):

It has the formula for compound interest - use A=55000 , n=4 , t=8, r=.05, and stick it into the formula:

OpenStudy (chris):

Then solve for P

OpenStudy (chris):

55000 = P*((1 + .05/4) ^ (4*8))

OpenStudy (chris):

once you multiply out the right hand side besides P, you'll have 55000 = P*(number from your calculator)

OpenStudy (chris):

then you can solve for P by doing P=55000/(number from calculator)

OpenStudy (chris):

55000/(1+.05/4)^(4*8) = 36,959.12